### v2

@@ 2.2 Selecting a number of checkpoints for the ensemble
We train 34 SFNO checkpoints from scratch, at which point we determined that 29 checkpoints adequately sample the ensemble spread as described below. We consider using different numbers of checkpoints in the size of the ensemble, from 4 checkpoints to 34 checkpoints, at intervals of 5 checkpoints. For each ensemble size, we conduct 100 bootstrap samples with replacement from the 34 checkpoints. Figure 3 shows the ensemble resulting spread obtained from these bootstrap samples. The ensemble spread is calculated as the global mean ensemble variance at each grid cell; it is calculated for a 120-hour lead time and averaged over forecasts initialized at 52 initialization dates (one initialization per week of 2018). We choose 120 hours because this timescale allows for synoptic-scale errors to grow, and given its importance for weather forecasting, we hope to represent model uncertainty for this time period as accurately as possible. Figure 3 shows that the ensemble spread asymptotes at approximately twenty-nine checkpoints. We conclude that twenty-nine checkpoints adequately sample the underlying population of all possible SFNO checkpoints with our selection of hyperparameters. In our ensemble results for the remainder of this paper, we use twenty-nine checkpoints. We open-source all 34 model checkpoints (each with 1.1 billion learned weights) as a resource to the community, to explore further the benefit of multiple SFNOs on forecasting atmospheric phenomena.
Figure 3: Ensemble spread from different numbers of checkpoints. Ensemble spread is calculated as the square root of time-mean, global-mean variance (Fortin et al., 2014). A correction factor of N-1 is applied to account for different ensemble sizes in the unbiased estimator of variance. At a lead time of five days, ensemble spread is averaged over forecasts from fifty-two initial conditions in the validation set (one per week starting 01-02-2018). For each number of SFNO checkpoints, 100 estimates of ensemble spread are obtained by taking 100 bootstrap random samples of the SFNO checkpoints. The box-and-whiskers visualize the distribution of these 100 trials.
@@ 2.3 Bred vectors with SFNO
    Generate spherical random noise correlated on 500 km length scales. Add this noise as a perturbation to 500 hPa geopotential at time $t_{-3}$
@@
    Generate a perturbed forecast using the perturbed input and a control forecast using the unperturbed input.
@@
The resulting perturbation is added to or subtracted from $t_{0}$. Using this perturbed initial condition, SFNO generates a 360-hour forecast, which serves a perturbed member in the ensemble.
@@
Figure 4: Diagram of generating bred vectors. This diagram details the process of generating bred vectors used for developing initial condition perturbations at $t_{0}$. First, using the input three time steps prior to $t_{0}$ (denoted $t_{-3}$), random noise is added to 500 hPa geopotential (z500). This noise respects spherical geometry and has a spatial correlation length scale of 500 km. With $t_{-3}$ as the initial condition, the perturbed forecast is subtracted from the control forecast. This difference is rescaled and used as a new perturbation, which is added to $t_{-2}$. This process is repeated for $t_{-1}$ and $t_{0}$. For each variable during every step of the breeding process, the amplitude of the perturbation is scaled to be 0.35 * the deterministic RMSE of SFNO at 48 hours
@@
The amplitude of the bred vectors is determined by the deterministic RMSE of SFNO at 48 hours, multiplied by a scaling factor of 0.35. At early lead times, the deterministic and ensemble mean RMSE of an ensemble forecast are similar. This factor is a tuning parameter. Since this parameter is less than 1, it reduces the perturbation amplitude. To satisfy criteria for statistical exchangeability, the ensemble spread should match its ensemble mean RMSE. Thus, we use the deterministic RMSE as a proxy for the desired spread level at early lead times. This approach provides a clear guide for the amplitude of each variable at each pressure level. Tuning these amplitudes across variables and levels would be challenging, since there are seventy-four different input variables.Figure 16 shows the actual amplitude for each of the seventy-four variables.
@@
We adopt two design choices from Toth and Kalnay (1997) and Toth and Kalnay (1993): centered perturbations and hemispheric-dependent amplitudes. For each learned bred vector, we both add and subtract it to the initial condition; this creates two separate perturbations (one positive and one negative). Centered perturbations improved the performance of the ensemble mean RMSE on the 2018 validation set (not shown). Additionally, we rescale the amplitudes separately for the Northern Hemisphere extratropics and the Southern Hemisphere extratropics. To prevent jump discontinuities in the perturbation amplitudes near 20 N and 20 S, a linearly interpolated rescaling factor is used in the tropics. Hemispheric rescaling prevents one hemisphere from dominating the perturbation amplitude. All perturbations are clipped to ensure that total column water vapor and specific humidity cannot be negative.
@@
In Step 2 of Figure 4, we add correlated spherical noise to 500 hPa geopotential (Z500). The noise has a correlation length scale of 500 km, and it has the same structure as noise of the Stochastic Perturbed Parameterized Tendency scheme used at ECMWF (Leutbecher and Palmer, 2008). We only add the initial noise to Z500, to avoid perturbing different fields in opposing and possibly contradictory directions. For instance, positively perturbing total column water vapor but negatively perturbing specific humidity at 1000, 925, and 850 hPa would likely be unphysical. Z500 is a natural choice of initial field to perturb because it is the steering flow in the extratropics. Because it is a smooth field on an isobaric surface, correlated spherical noise is an appropriately structured additive perturbation. On the other hand, correlated spherical noise would not serve well as an additive perturbation to surface fields, which have sharp discontinuities due to orography and land sea contrasts. We design the bred vectors with the goal of keeping the perturbed input as close to the training dataset as possible. We minimize the extent of directly prescribed perturbations, and the majority of the perturbation structure is generated from the breeding process with SFNO itself. To start the breeding cycle, the initial perturbation is applied to Z500, but for all subsequent cycles, all 74 input variables are perturbed. In this manner, we develop a mutually consistent way of perturbing all input channels.
@@
We test our bred vectors by evaluating spread-error performance on the validation year: 2018. Figure 5 visualizes sample bred vectors for various input fields and channels. These perturbations contain some desirable qualities. First, they contain a land-sea contrast for surface fields such as 10m wind speed and 2m temperature. For these surface fields, perturbations have distinct amplitudes and spatial scales over the land and ocean. The specific humidity perturbations are stronger in the tropics than at the poles, in line with a strong equator-to-pole moisture gradient. This is a benefit of using bred vectors, compared to perturbing multiple input variables simply with simple spherical noise, as in GraphCast-Perturbed (Price et al., 2023).
@@
We initially presented bred vectors and multiple checkpoints in Collins et al. (2024). Concurrently, Baño-Medina et al. (2024) also released a preprint using bred vectors and multiple trained models. The results in Baño-Medina et al. (2024) serve as excellent independent validation of bred vectors and multiple checkpoints. They validate their method from Jan 10 to Feb 28 (with 50 forecast initial dates), and they show promising results, particularly at certain latitudes and land regions. We comprehensively show that SFNO-BVMC is competitive with IFS on global mean quantities using forecasts from a full year (732 forecast initial dates for 2020 and 92 for summer 2023). We further validate our ensemble with a unique pipeline for extreme diagnostics and spectral diagnostics of each ensemble member and the ensemble mean. While their method uses Adaptive Fourier Neural Operators (AFNO) (Pathak et al., 2022), we use SFNO, a successor to AFNO that is more stable and has better skill. We train all 29 SFNOs from scratch, whereas they sample multiple models from 3 training runs. To compare methodologies with Figure 2 in Baño-Medina et al. (2024), we present a diagram of how we generated bred vectors in SFNO-BVMC. The boxed quantities in Figure 4 represent the unique methodological details of our approach. We add spherical initial noise to Z500 (compared to Gaussian noise), start the breeding cycle 3 timesteps before the initialization date (compared to Jan 1, 2018), and use the deterministic RMSE as the bred vector amplitude. In Part II, we assess the forecasts from bred vectors and multiple checkpoints at scale, with a larger ensemble.
@@ 2.4 Contributions of bred vectors and multiple checkpoints to the ensemble calibration
In SFNO-BVMC, the bred vectors and multiple trained model checkpoints both contribute to ensemble spread and calibration. Bred vectors are a flow-dependent initial condition perturbation: they are calculated independently for each checkpoint, and they use the preceding three time steps to generate the perturbation according to the current flow in the atmosphere. At longer lead times, when there is less dependence on the initial conditions, multi-checkpointing causes the spread-error ratio to approach 1; this is consistent with our expectations from the grand ensemble in Weyn et al. (2021). In Figure 6, we show the spread-error ratios from three different ensembles: Figure 6a has only 29 checkpoints and no bred vectors, Figure 6b has 1 checkpoint and 29 bred vectors (each added to and subtracted the initial condition), and Figure 6c has 29 checkpoints and 1 bred vector (added to and subtracted from the initial condition). Figure 6a has 29 ensemble members, while Figure 6s b and c have 58 members ensembles. As a model perturbation, multi-checkpointing does not represent the uncertainty arising from an imperfect initial condition. Therefore, the multi-checkpoint ensemble is underdispersive at early lead times. On the other hand, the ensemble composed only of bred vectors is underdispersive on synoptic time scales (3-5 days) when representing model uncertainty also becomes important for obtaining good calibration (Palmer, 2018).
@@ 3 Ensemble Diagnostics
Ultimately, with SFNO-BVMC, we hope to analyze the statistics of LLHIs. This requires a calibrated ensemble with reliable probabilistic forecasts. SFNO-BVMC is a novel way to create ensemble forecasts from deterministic ML models. Therefore, in the following section, we present a diagnostics pipeline to evaluate the SFNO-BVMC ensemble and compare it to the IFS ensemble. We first evaluate SFNO-BVMC using diagnostics that evaluate overall performance. Next, we assess SFNO-BVMC’s control, perturbed, and ensemble mean spectra. Finally, we present diagnostics specifically focused on extreme weather forecasts. We open-source the code for these diagnostics (see Data Availability section), and we hope that it can be used to guide future ML model development. For a fair comparison for all diagnostics, we validate IFS against ECMWF’s operational analysis and SFNO-BVMC against ERA5. All diagnostics show SFNO-BVMC results with 58 members, and IFS ENS results with 50 members. SFNO-BVMC has 58 members: 29 checkpoints and 1 bred vector per checkpoint (added to and subtracted from the initial condition). Because of the use of 29 checkpoints and centered bred vector perturbations, SFNO-BVMC cannot be evaluated with an ensemble size smaller than fifty-eight members. However, while there are unbiased versions of the metrics, the difference in the metrics due to different ensemble size would be sufficiently small that the diagnostics still allow for fair evaluation.
@@ 3.1 Mean Diagnostics
Figure 7 compares the global mean CRPS of SFNO-BVMC to that of IFS ENS on five different variables. On 850 hPa temperature, 2m temperature, 850 hPa specific humidity, and 500 hPa geopotential, SFNO-BVMC lags approximately 18 hours behind IFS ENS, though their performance is comparable. SFNO-BVMC does match IFS ENS on the 10m zonal (u component) wind.
@@
Second, an essential requirement for an ensemble weather forecast is that the ensemble spread must match its skill (Fortin et al., 2014); the spread-error ratio should be 1. This result is derived statistically based on the idea of exchangeability between ensemble members: each ensemble member should be statistically indistinguishable from each other and from the forecasts (Fortin et al., 2014; Palmer et al., 2006). The spread is the square root of the global-mean ensemble variance. Similarly, the error is the square root of the global-mean ensemble MSE. See Section B for a detailed description of calculating the spread and error across multiple forecasts initialized on different initial times. Figure 8 demonstrates that SFNO-BVMC obtains spread-error ratios that approach 1, and it has comparable performance to IFS ENS. At early lead times, SFNO-BVMC is underdispersive for all variables except Z500, but the spread skill ratio approaches 1 for longer lead times.
@@
Finally, we evaluate the ensemble mean RMSE of SFNO-BVMC and IFS ENS (Figure 9). Their scores are comparable, and both models have an ensemble mean RMSE that converges to climatology at 360 hours (14 days).
@@
On these aggregate metrics, SFNO-BVMC is often eighteen hours behind IFS ENS, so its performance is slightly worse but still comparable. Through large SFNOs with a high-resolution, expressive internal state, bred vectors, and multi-checkpointing, this ensemble has significantly improved calibration, compared to previous work using lagged ensembles (Brenowitz et al., 2024). It serves as a benchmark for the calibration potential for deterministic ML models, and it can be compared to recent models which optimize for an ensemble objective. However, SFNO-BVMC is orders of magnitude less computationally expensive, so it uniquely enables the creation of huge ensembles of forecasts. These allow for unprecedented sampling of internal variability and an analysis of extreme statistics, as presented in Part II of this paper. Additionally, while IFS ENS has been an established weather forecasting model for decades, SFNO is still a new architecture. Improving the skill of the SFNO architecture itself is an important area of future research.

### v3

@@ 2.2 Selecting a number of checkpoints for the ensemble
We determine that 29 checkpoints adequately sample the ensemble spread. We experiment with using different numbers of checkpoints in the size of the ensemble, from 4 checkpoints to 34 checkpoints, at intervals of 5 checkpoints. For each ensemble size, we conduct 100 bootstrap samples with replacement from the 34 checkpoints. Figure 3 shows the resulting ensemble spread obtained from these bootstrap samples. The ensemble spread is calculated as the global mean ensemble variance at each grid cell; it is calculated for a 120-hour lead time and averaged over forecasts initialized at 52 initialization dates (one initialization per week of 2018). We choose 120 hours because this timescale allows for synoptic-scale errors to grow, and given its importance for weather forecasting, we hope to represent model uncertainty for this time period as accurately as possible. Figure 3 shows that the ensemble spread asymptotes at approximately twenty-nine checkpoints. We conclude that twenty-nine checkpoints adequately sample the underlying population of all possible SFNO checkpoints with our selection of hyperparameters. In our ensemble results for the remainder of this paper, we use twenty-nine checkpoints. We open-source all 34 model checkpoints (each with 1.1 billion learned weights) as a resource to the community, to explore the benefit of multiple SFNOs on forecasting atmospheric phenomena.
Figure 3: Ensemble spread from different numbers of checkpoints. Ensemble spread is calculated as the square root of time-mean, global-mean variance (Fortin et al., 2014). A correction factor of N-1 is applied to account for different ensemble sizes in the unbiased estimator of variance. At a lead time of five days, ensemble spread is averaged over forecasts from fifty-two initial conditions in the validation set (one per week starting 01-02-2018). Ensemble spread is shown for total column water vapor (left), 10m wind speed (middle), and 2m temperature (right). For each number of SFNO checkpoints, 200 estimates of ensemble spread are obtained by taking 100 bootstrap random samples of the SFNO checkpoints. The box-and-whiskers visualize the distribution of these 200 trials: the middle of the box is the median, the ends of the box are the first and third quartile of the data, and the ends of the box are correspond to the minimum and maximum.
@@ 2.3 Bred vectors with SFNO
    Generate spherical random noise correlated on 500 km length scales. Add this noise as a perturbation to 500 hPa geopotential at time $t_{-2}$
@@
    Generate a perturbed forecast (with the perturbed input) and a control forecast (with the unperturbed input).
@@
The resulting perturbation can be added to or subtracted from $t_{0}$. Using this perturbed initial condition, SFNO generates a 360-hour forecast, which serves a perturbed member in the ensemble.
@@
Figure 4: Diagram of generating bred vectors. This diagram details the process of generating bred vectors used for developing initial condition perturbations at $t_{0}$. First, using the input three time steps before $t_{0}$ (denoted $t_{-2}$), random noise is added to 500 hPa geopotential (z500). This noise respects spherical geometry and has a spatial correlation length scale of 500 km. With $t_{-2}$ as the initial condition, the perturbed forecast is subtracted from the control forecast. This difference is rescaled and used as a new perturbation, which is added to $t_{-1}$. This process is repeated for $t_{0}$. For each variable during every step of the breeding process, the amplitude of the perturbation is scaled to be 0.35 * the deterministic RMSE of SFNO at 48 hours
@@
The amplitude of the bred vectors is determined by the deterministic RMSE of SFNO at 48 hours, multiplied by a scaling factor of 0.35. This factor is a tuning parameter. Since this parameter is less than 1, it reduces the perturbation amplitude. At early lead times, the deterministic and ensemble mean RMSE of an ensemble forecast are similar. To satisfy criteria for statistical exchangeability, the ensemble spread should match its ensemble mean RMSE. We use the deterministic RMSE (with a tuning parameter) as a proxy for the desired spread level at early lead times. This approach provides a clear guide for the amplitude of each variable at each pressure level. Manually tuning these amplitudes across variables and levels would be challenging, since there are seventy-four different input variables. Figure 17 shows the actual amplitude for each of the seventy-four variables.
@@
We adopt two design choices from Toth and Kalnay (1997) and Toth and Kalnay (1993): centered perturbations and hemispheric-dependent amplitudes. For each learned bred vector, we both add it to and subtract it from the initial condition; this creates two separate perturbations. Centered perturbations improved the performance of the ensemble mean RMSE on the 2018 validation set. Additionally, we rescale the amplitudes separately for the Northern Hemisphere extratropics and the Southern Hemisphere extratropics. To prevent jump discontinuities in the perturbation amplitudes near 20 N and 20 S, a linearly interpolated rescaling factor is used in the tropics. Hemispheric rescaling prevents one hemisphere from dominating the perturbation amplitude. All perturbations are clipped to ensure that total column water vapor and specific humidity cannot be negative. See Appendix Section E for a note about our implementation of bred vectors.
@@
In Step 2 of Figure 4, we add correlated spherical noise to 500 hPa geopotential (Z500). The noise has a correlation length scale of 500 km, and it has the same structure as noise of the Stochastic Perturbed Parameterized Tendency scheme used at ECMWF (Leutbecher and Palmer, 2008). We only add the initial noise to Z500, to avoid perturbing different fields in opposing and possibly contradictory directions. For instance, positively perturbing total column water vapor but negatively perturbing specific humidity on the lower pressure levels would likely be unphysical. Z500 is a natural choice of initial field to perturb because it is the steering flow in the extratropics. Since it is a smooth field on an isobaric surface, correlated spherical noise is an appropriately structured additive perturbation. On the other hand, correlated spherical noise would not serve well as an additive perturbation to surface fields, which have sharp discontinuities due to orography and land sea contrasts. We design the bred vectors with the goal of keeping the perturbed input as close to the training dataset as possible. We minimize the extent of directly prescribed perturbations, and the majority of the perturbation structure is generated from the breeding process with SFNO itself. To start the breeding cycle, the initial perturbation is applied to Z500, but for all subsequent cycles, all 74 input variables are perturbed. In this manner, we develop a mutually consistent way of perturbing all input channels.
@@
We test our bred vectors by evaluating spread-error performance on the validation year: 2018. Figure 5 visualizes sample bred vectors for various input fields and channels. These perturbations contain some desirable qualities. First, they contain a land-sea contrast for surface fields such as 10m wind speed and 2m temperature. In this example, the 2m temperature perturbation has an amplitude of 0.56 K over land and 0.27 K over the ocean, and the 10m wind speed perturbation has an amplitude of 0.45 m/s over land and 0.66 m/s over the ocean. The specific humidity perturbations are stronger in the tropics than at the poles, in line with the equator-to-pole moisture gradient. These physical qualities of bred vectors are a benefit of using bred vectors over simple spherical noise, as in GraphCast-Perturbed (Price et al., 2023), or Perlin noise (Bi et al., 2023).
@@
We initially presented bred vectors and multiple checkpoints in Collins et al. (2024). Concurrently, Baño-Medina et al. (2024) released a preprint using bred vectors and multiple trained models. The results in Baño-Medina et al. (2024) serve as excellent independent validation of bred vectors and multiple checkpoints. They validate their method from Jan 10 to Feb 28 (with 50 forecast initial dates), and they show promising results, particularly at certain latitudes and land regions. We comprehensively show that SFNO-BVMC is competitive with IFS on global mean quantities using forecasts from a full year (732 forecast initial dates for 2020 and 92 for summer 2023). We further validate our ensemble with a unique pipeline for extreme diagnostics and spectral diagnostics of each ensemble member and the ensemble mean. While their method uses Adaptive Fourier Neural Operators (AFNO) (Pathak et al., 2022), we use SFNO, a successor to AFNO that is more stable and has better skill. We train all 29 SFNOs from scratch, whereas they sample multiple models from 3 training runs. To compare methodologies with Figure 2 in Baño-Medina et al. (2024), we present a diagram of how we generated bred vectors in SFNO-BVMC. The boxed quantities in Figure 4 represent the unique methodological details of our approach. We add spherical initial noise to Z500 (compared to Gaussian noise), start the breeding cycle 3 timesteps before the initialization date (compared to Jan 1, 2018), and use the deterministic RMSE as the bred vector amplitude. In Part II, we assess the forecasts from bred vectors and multiple checkpoints at scale, with a significantly larger ensemble than in Baño-Medina et al. (2024).
@@ 2.4 Contributions of bred vectors and multiple checkpoints to the ensemble calibration
In SFNO-BVMC, the bred vectors and multiple trained model checkpoints both contribute to ensemble spread and calibration. Bred vectors are a flow-dependent initial condition perturbation: they are calculated independently for each checkpoint, and they use the preceding two time steps to generate the perturbation based on the current flow in the atmosphere. At longer lead times, when there is less dependence on the initial conditions, multi-checkpointing causes the spread-error ratio to approach 1; this is consistent with our expectations from the ensemble in Weyn et al. (2021). In Figure 6, we show the spread-error ratios from three different ensembles: Figure 6a has only 29 checkpoints and no bred vectors, Figure 6b has 1 checkpoint and 29 bred vectors (each added to and subtracted the initial condition), and Figure 6c has 29 checkpoints and 1 bred vector (added to and subtracted from the initial condition). Figure 6a has 29 ensemble members, while Figures 6b and c have 58-member ensembles. As a model perturbation, multi-checkpointing does not represent the uncertainty arising from an imperfect initial condition. Therefore, the multi-checkpoint ensemble is underdispersive at early lead times. On the other hand, the ensemble composed only of bred vectors is underdispersive on synoptic time scales (3-5 days) when representing model uncertainty also becomes important for obtaining good calibration.
@@ 3 Ensemble Diagnostics
Ultimately, with SFNO-BVMC, we hope to study LLHIs. This requires a calibrated ensemble with reliable probabilistic forecasts. SFNO-BVMC is a novel way to create ensemble forecasts from deterministic ML models. Therefore, in the following section, we present a diagnostics pipeline to evaluate the SFNO-BVMC ensemble and compare it to the IFS ensemble. We first evaluate SFNO-BVMC using diagnostics that evaluate overall performance. Next, we assess SFNO-BVMC’s control, perturbed, and ensemble mean spectra. Finally, we present diagnostics specifically focused on extreme weather forecasts. We open-source the code for these diagnostics (see Data Availability section), and we hope that it can be used to guide future ML model development. For a fair comparison for all diagnostics, we validate IFS against ECMWF’s operational analysis and SFNO-BVMC against ERA5. IFS is initialized with this operational analysis, not the ERA5 reanalysis, so it has a different verification dataset. All diagnostics show SFNO-BVMC scores with 58 members, and IFS ENS scores with 50 members. SFNO-BVMC has 58 members: 29 checkpoints and 1 bred vector per checkpoint (added to and subtracted from the initial condition). Because of the use of 29 checkpoints and centered bred vector perturbations, SFNO-BVMC cannot be evaluated with an ensemble size smaller than fifty-eight members. While there are versions of the metrics that are corrected for ensemble size, the difference in the metrics due to different ensemble size would be sufficiently small that the diagnostics still allow for comparison between the 50-member IFS and 58-member SFNO-BVMC.
@@ 3.1 Mean Diagnostics
Figure 7 compares the global mean CRPS of SFNO-BVMC to that of IFS ENS on five different variables. On 850 hPa temperature, 2m temperature, 850 hPa specific humidity, and 500 hPa geopotential, SFNO-BVMC lags approximately 12–18 hours behind IFS ENS. SFNO-BVMC does match IFS ENS on the 10m zonal (u component) wind.
@@
Second, an essential requirement for an ensemble weather forecast is that the ensemble spread must match its skill (Fortin et al., 2014); the spread-error ratio should be 1. This result is derived statistically based on the idea of exchangeability between ensemble members: each ensemble member should be statistically indistinguishable from each other and from the forecasts (Fortin et al., 2014; Palmer et al., 2006). The spread is the square root of the global-mean ensemble variance. Similarly, the error is the square root of the global-mean ensemble MSE. See Section C for a detailed description of calculating the spread and error across multiple forecasts initialized on different initial times. Figure 8 demonstrates that SFNO-BVMC obtains spread-error ratios that approach 1, and it has comparable performance to IFS ENS. At early lead times, SFNO-BVMC is underdispersive for all variables except Z500, but the spread skill ratio approaches 1 for longer lead times.
@@
Finally, we evaluate the ensemble mean RMSE of SFNO-BVMC and IFS ENS (Figure 9). Their scores are comparable, with SFNO-BVMC lagging close behind the IFS ensemble mean, and both models have an ensemble mean RMSE that converges to climatology at 360 hours (14 days).
@@
Through large SFNOs with a high-resolution, expressive internal state, bred vectors, and multi-checkpointing, this ensemble has significantly improved calibration, compared to previous work using lagged ensembles (Brenowitz et al., 2024). It serves as a benchmark for the calibration potential for deterministic ML models, and it can be compared to recent models which optimize for an ensemble objective. While IFS ENS has been an established weather forecasting model for decades, SFNO is still a new architecture. Improving the skill of the SFNO architecture itself is an important area of future research. However, in this manuscript, our main goal is not primarily to create the most skillful weather forecasting model; rather, we hope to explore huge ensembles and low-likelihood events at the tail of the ensemble forecast distribution. SFNO-BVMC is less computationally expensive than IFS, so it uniquely enables the creation of huge ensembles. These allow for unprecedented sampling of internal variability and an analysis of extreme statistics, as presented in Part II of this paper.